a small oxygen cylinder at 300k and 10 bar is being filled from a larger tank which is at 300k and 100 bar, constant throughout the process. calculate the final temperature in the small cylinder when the pressure reaches 80 bar a) assuming that o2 is an ideal gas with cv=5r/2 and b)using accurate thermodynamic data for o2
1
Expert's answer
2012-10-01T10:11:45-0400
a) Cv=x*80
Cv=5r/2
5r=x*160
r=160*37.5/5=1200

b)
300*10=x*80
x=37.5